OpenStudy (anonymous):

find the three consecutive even integers such that the first times the third is 2 greater than 5 times the second?

terenzreignz (terenzreignz):

http://openstudy.com/study#/updates/4feae23ae4b0bbec5cfaa74d

Parth (parthkohli):

x, x + 2, x + 4 can be the three even integers. \( \color{Black}{\Rightarrow x(x + 4) = 2 + 5(x + 2)}\) \( \color{Black}{\Rightarrow x^2 + 4x =2 + 5x + 10}\) \( \color{Black}{\Rightarrow x^2 + 4x = 12 + 5x }\) \( \color{Black}{\Rightarrow x^2 - x - 12 = 0}\) Solve this quadratic equation.

OpenStudy (amistre64):

2x is an even integer; by default tho

OpenStudy (amistre64):

x, x+2, x+4 are just 3 consecutive even or odd integers i think

OpenStudy (amistre64):

2x, 2x+1, 2x+2 would be 3 consecutive even integers by default

Parth (parthkohli):

Uh oh, but it gets you even integers actually

terenzreignz (terenzreignz):

@amistre64 OH NO 2x + 1 is ODD if 2x is even O.O

Parth (parthkohli):

Well actually solve it and then see the even solution.

OpenStudy (amistre64):

hmmm, math has changed since the 60s it appears :) answers are a consortium these days lol

terenzreignz (terenzreignz):

All I meant is that any even number, if you add one to it, it'll be odd, right? :)

OpenStudy (amistre64):

that is correct, i had half an idea and the other side of my brain forgot to chime in ;) good catch tho

terenzreignz (terenzreignz):

haha, me and my anxiety attacks :D

OpenStudy (amistre64):

this seems to be a popular question, as if it was on some sort of quiz or exam ...

OpenStudy (amistre64):

last time we had a bout of same question-itis ... a teacher had told their students to go ask it on openstudy

OpenStudy (anonymous):

amistre64,terenzreignz, parthkholi thanks :)